OpenStudy (anonymous):

What is the slope-intercept form of the equation y - 7 = - 5/2(x + 4)? o y - 2 = -5/2(x + 2) o y + 7 = -x + 5/2 o y = - 4/7 + 2 o y = - 5/2x - 3

OpenStudy (nubeer):

y=mx + c this is slope-intercept form.. just see which of your equation looks like this..

OpenStudy (anonymous):

Okay, I think it would be the second option then; y = 2/5x + 2.

OpenStudy (anonymous):

Oh sorry, I'm wrong. I put the wrong choices by accident.

OpenStudy (nubeer):

hmm ya i noticed your question gives a different answer

OpenStudy (anonymous):

I fixed the original question options.

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Okay, I think for this particular problem (in the original question), the answer is the third choice, since it is the only one in slope-intercept form and has a positive b. Am I correct?

OpenStudy (nubeer):

nop.. its not 3rd check again what slope u have in question. and match it with the options.

OpenStudy (anonymous):

not c, slope is different than what's given

OpenStudy (anonymous):

Okay, well, I first need to times /5/2(x + 4) right? I need to actually work out the problem rather than see which one it matches with?

OpenStudy (anonymous):

I meant -5/2

OpenStudy (anonymous):

I see now why it isn't the third one.

OpenStudy (anonymous):

I would say the fourth one. -7 + 4 = -3 for b. Correct?